# Next-to-leading-order QCD corrections to nucleon Dirac form factors

### v2

@@ Abstract
$+e_{d}\,d^{ijkl}_{0}\left[1+{\alpha_{s}\over\pi}(d^{ijkl}_{1}L_{\mu}+d^{ijkl}_{2})\right]+{\cal O}(\alpha_{s}^{2})\Big{\}},$
with $L_{\mu}\equiv\ln\mu^{2}/Q^{2}$. For simplicity we have set $\mu_{R}=\mu_{F}=\mu$. We find $c^{ijkl}_{1}=d^{ijkl}_{1}$, and ${\cal H}^{ijkl}={\cal H}^{klij}$ originating from $T_{H}(x,y)=T_{H}(y,x)$. For a given $ijkl$, the coefficients $c^{ijkl}_{012}$ and $d^{ijkl}_{012}$ can be analytically inferred. For concreteness, we just list their values for the case of $0000$:
${\cal H}^{0000}$ $=$ $\frac{1600\alpha_{s}^{2}\pi^{2}}{3}\Big{\{}e_{u}\left[1+\frac{\alpha_{s}}{\pi}\left(\frac{29}{6}L_{\mu}-92\zeta_{5}\right.\right.$ (23)
$\left.\left.+\frac{3719}{30}\zeta_{3}-\frac{1897}{60}\right)\right]+2e_{d}\left[1+\frac{\alpha_{s}}{\pi}\left(\frac{29}{6}L_{\mu}+\frac{3}{2}\zeta_{5}\right.\right.$
$\left.\left.+\frac{209}{120}\zeta_{3}+\frac{1741}{120}\right)\right]+{\cal O}(\alpha_{s}^{2})\Big{\}},$
${ijkl}$ $c^{ijkl}_{0}$ $c^{ijkl}_{2}$ $d^{ijkl}_{0}$ $d^{ijkl}_{2}$ $c^{ijkl}_{1}=d^{ijkl}_{1}$
$0000$ 1.0000 22.001 2.0000 18.157 4.8333
$1000$ 8.1667 25.030 -8.1667 25.030 5.3889
$1010$ 168.78 29.453 174.22 28.937 5.9444
$1100$ 3.5000 29.699 -3.5000 18.566 5.5000
$1110$ 38.111 30.441 -38.111 30.441 6.0556
$1111$ 59.889 27.617 21.778 26.486 6.1667
$2000$ 21.350 26.952 8.0500 25.082 5.7222
$2010$ 89.833 32.057 -89.833 32.057 6.2778
$2011$ 93.100 32.942 9.8000 38.805 6.3889
$2020$ 234.71 35.077 54.880 34.146 6.6111
$2100$ 8.7500 28.194 -8.7500 28.194 5.9444
$2110$ 159.25 33.344 159.25 33.569 6.5000
$2111$ 44.917 33.321 -44.917 33.321 6.6111
$2120$ 84.525 36.543 -84.525 36.543 6.8333
$2121$ 159.25 38.254 159.25 38.232 7.0556
$2200$ 3.2000 27.963 0.10000 47.505 6.0000
$2210$ 12.250 32.193 -12.250 32.193 6.5556
$2211$ 7.1167 40.486 2.6833 35.971 6.6667
$2220$ 30.345 36.568 4.9350 37.529 6.8889
$2221$ 7.1750 39.572 -7.1750 39.572 7.1111
$2222$ 7.7900 35.348 1.7200 35.302 7.1667
@@
Conducting the four-fold convolution integrations to accurately inferring ${\cal H}^{ijkl}$ turns out to be a great challenge, largely due to severe numerical instability caused by many fake singularities in $T_{H}^{(1)}(x,y)$, which are located either at the boundaries or in the interior of the integration domain. We conduct the four-fold integration from two different routes: one is to utilize the numerical integrator Cubature cubature implementing multiple-precision floating-point around 100 digits, and the other is to accomplish the integration analytically with the assistance of the package PolyLogTools Duhr:2019tlz . The perfect agreement is found between the analytic and numerical approaches. In Table 1 we enumerate the values for $c_{i}$ and $d_{i}$ associated with a few lowest-lying $ijkl$  [^7] We notice that, a PhD thesis by Knödlseder in 2015 was dedicated to calculating the NLO QCD correction to nucleon Dirac form factor at leading twist Knodlseder:2015vmu . Though the analytic result of the one-loop amplitude was obtained in Knodlseder:2015vmu , the IR-finite hard-scattering kernel $T^{(1)}(x,y)$ has not been given in a pointwise manner. More importantly, since the author failed to accomplish the four-fold convolution integrations in a numerically stable fashion, no attempt of confronting the NLO pQCD predictions with the measured nucleon form factor data was made in Knodlseder:2015vmu ..
Equipped with Table 1, Equation (20) constitutes the master formula for yielding phenomenological predictions for proton Dirac form factor through the one-loop accuracy. In the isospin symmetric limit, the neutron Dirac form factor $F_{1}^{n}$ can be obtained from (20) by making the replacement $e_{u}\leftrightarrow e_{d}$ in (22). It is also straightforward to adapt this master formula from the space-like region to the time-like one, provided that one makes the replacement $L_{\mu}\to L_{\mu}+i\pi$ in (22), with $Q^{2}$ now indicating the squared invariant mass of the nucleon-antinucleon pair.
Phenomenological exploration. It is now the time to confront our state-of-the-art pQCD predictions with the available data. We take the large-$Q^{2}$ data points for spacelike proton Dirac form factor from SLAC 1986 Coward:1967au and SLAC 1993 Sill:1992qw   [^8] Note it is the magnetic Sachs form factor of nucleon, $G_{\rm M}(Q^{2})$, that is directly accessible in $ep$ scattering experiments. At large $Q^{2}$, it nevertheless becomes legitimate to assume $F_{1}(Q^{2})\approx G_{\rm M}(Q^{2})$.. We take the data points for time-like proton Dirac form factor from BES BES:2005lpy , CLEO CLEO:2005tiu , NU Seth:2012nn , E835 E835:1999mlt , E760 E760:1992rvj , LA ISR BaBar:2013ves and SA ISR BaBar:2013ukx . We discard many irrelevant data at small momentum transfer. In contrast, the quality of data and the covered range of $Q^{2}$ for neutron EMFFs is much less competitive with respect to that for proton. The data points for spacelike neutron Dirac form factor are taken from ROCK Rock:1982gf , Lung Lung:1992bu , while those for timelike neutron Dirac form factor are taken from FENICE Antonelli:1998fv and BES3 BESIII:2021tbq .
@@
As the key nonperturative input, our present knowledge about the nucleon LCDAs is far from satisfactory and confirmative. During the past decades, The low-lying expansion coefficients of nucleon LCDAs have been extensively investigated Chernyak:1984bm ; Chernyak:1987nv ; King:1986wi ; Stefanis:1992nw ; Braun:2006hz ; Anikin:2013aka ; Braun:2014wpa ; RQCD:2019hps . In our numerical analysis, we have taken eight different ansatzs of nucleon LCDAs as input. Six of them are obtained from the QCD sum rules, referred to as CZ Chernyak:1984bm , COZ Chernyak:1987nu , KS King:1986wi , SB Stefanis:1992nw , BLW Braun:2006hz and ABO Anikin:2013aka , respectively. Two of them are provided by the recent lattice QCD simulations, abbreviated as LAT 14 Braun:2014wpa and RQCD RQCD:2019hps , respectively.
The predicted values of the expansion coefficients of nucleon are usually given at the initial scale $\mu_{0}=1$ or 2 GeV, which are scattered in a wide range in literature. The one-loop renormalization group equations are invoked to evolve $f_{\rm N}$ and $a_{nk}(\mu)$ from $\mu_{0}$ to another scale, viz. $f_{\rm N}(\mu)=f_{\rm N}(\mu_{0})\left[\alpha_{s}(\mu)/\alpha_{s}(\mu_{0})\right]^{2/(3\beta_{0})}$ and $a_{ij}(\mu)=a_{ij}(\mu_{0})\left[\alpha_{s}(\mu)/\alpha_{s}(\mu_{0})\right]^{\gamma_{ij}/\beta_{0}}$. $\beta_{0}=11-{2\over 3}n_{f}$ with $n_{f}=3$ signifies the one-loop coefficient of the QCD $\beta$ function, while $\gamma_{ij}$ is the anomalous dimension associated with each expansion coefficient, whose explicit form can be found in Braun:2014wpa . The package FAPT Bakulev:2012sm is utilized to evaluate the running QCD coupling constant to two-loop accuracy.
@@
It has been widely known that the LO predictions in pQCD for proton Dirac form factor are generally much smaller than the large-$Q^{2}$ data, for virtually all the popular models of nucleon LCDAs Stefanis:1997zyh . To resolve this discrepancy, it has been proposed Chernyak:1984bm ; Gari:1986ue ; Stefanis:1987vr ; Gari:1986dr ; Ji:1986uh ; Stefanis:1992nw that the strong coupling constants should be evaluated at the exact gluon virtualities in each diagram, so that the LO prediction can be significantly enhanced to account for the data. The $\alpha_{s}$ in this recipe might be associated with a scale of order $\Lambda_{\rm QCD}$, evaluation of which requires some ad hoc treatment. Moreover, it is difficult to adapt this recipe to the NLO case.
As is evident from Fig. 2 through Fig. 5, the magnitudes of the NLO QCD correction are positive and significant. After including the NLO QCD correction, the pQCD predictions, within errors, can decently account for the large-$Q^{2}$ data for both proton and neutron’s Dirac form factors, in both spacelike and timelike regions, once we adopt the CZ Chernyak:1984bm , COZ Chernyak:1987nu , KS King:1986wi , SB Stefanis:1992nw parametrizations of nucleon LCDAs. This phenomenological success of NLO prediction implies that the frozen $\alpha_{s}$ trick is unnecessary and should be abandoned.
@@
In sharp contrast, due to very small expansion coefficients predicted by BLW Braun:2006hz , ABO Anikin:2013aka , LAT14 Braun:2014wpa and RQCD RQCD:2019hps , the resulting pQCD predictions are far below the experimental data. From Fig. 2 through Fig. 5, one clearly sees that, even after including NLO corrections, the pQCD predictions based on these models are still order-of-magnitude smaller than the measured Dirac form factor of proton and neutron.
@@
Since lattice QCD is universally believed to be the only model-independent nonperturbative approach to unravel the hadronic structure, the sheer failure of the parameterized nucleon LCDAs provided by LAT14 and RQCD when confronting with the data is rather disquieting. This nuisance is apparently due to too small values of the Appel moments determined by the lattice method. To settle this alarming discrepancy, it seems imperative to call for another independent lattice study based on a diametrically different algorithm. The large momentum effective theory (LaMET) Ji:2013dva is a viable option, which allows one to ascertain hadron’s parton distributions directly in the $x$ space on Euclidean lattice (for a recent review on LaMET, see Ji:2020ect ). Some theoretical investigations have recently been launched to infer light octet baryons’ LCDAs within the LaMET framework Deng:2023csv ; Han:2024ucv .
@@
Summary and Outlook. In this work we report the first successful calculation of the NLO QCD corrections to the nucleon Dirac form factors. The celebrated collinear factorization formalism has been explicitly verified to hold true at one-loop order for hard-exclusive process involving baryons. Somewhat surprisingly, it takes nearly half a century to advance the LO prediction to the NLO level for such a basic observable related to nucleon internal structure.
It has long been known that the LO prediction in pQCD severely underestimates the measured nucleon Dirac form factor data. Historically some ad hoc frozen strong coupling recipe were advocated to reconcile this discrepancy. Remarkably, the effect of NLO perturbative corrections turns to be positive and significant. After including the NLO QCD correction, the pQCD predictions can decently describe the large-$Q^{2}$ data for both proton and neutron’s Dirac form factors, in both spacelike and timelike regions, provided that the parameterized forms of the nucleon LCDAs are taken from a class of QCD sum rules-based models, which renders the aforementioned frozen $\alpha_{s}$ trick unnecessary and obsolete. On the other hand, some other models and lattice QCD studies predict too small expansion coefficients of nucleon LCDAs, so that the resulting pQCD predictions, even after including NLO corrections, are still orders-of-magnitude smaller than the measured nucleon form factors.
@@
Acknowledgments. We are grateful to Deshan Yang for valuable discussion. The work of L. B. C. is supported in part by NNSFC under Grant No. 12175048. The work of W. C. is supported by Guangdong Major Project of Basic and Applied Basic Research (No. 2020B0301030008). The work of F. F. is supported in part by NNSFC under Grant No. 12275353. The work of S. H. and Y. J. is supported in part by NNSFC under Grant No. 11925506.
Note added. After the first version of this paper was submitted to arXiv, there has appeared a similar work by Huang, Shi, Wang and Zhao [arXiv:2407.18724[hep-ph]], who have also computed the NLO QCD correction to nucleon Dirac form factor in the spacelike domain, yet within a specific evanescent-to-physical operator mixing scheme.
Appendix A One-loop renormalization factor of non-local three-quark operators in KM scheme
To the best of our knowledge, the explicit one-loop expression of the renormalization factor ${\cal Z}$ in (8), which is associated with the nonlocal three-quark opeartors with light-like separations, has never been given in literature. For reader’s convenience, here we explicit present its closed form in KM scheme Krankl:2011gch , accurate through the one-loop accuracy:

### v3

@@ Abstract
$\hskip-19.91684pt+e_{d}\,d^{ijkl}_{0}\left[1+{\alpha_{s}\over\pi}(d^{ijkl}_{1}L_{\mu}+d^{ijkl}_{2})\right]+{\cal O}(\alpha_{s}^{2})\Big{\}},$
with $L_{\mu}\equiv\ln\mu^{2}/Q^{2}$. For simplicity we have set $\mu_{R}=\mu_{F}=\mu$. Note ${\cal H}^{ijkl}={\cal H}^{klij}$ and $c^{ijkl}_{1}=d^{ijkl}_{1}$. For a given ijklijkl, the coefficients $c^{ijkl}_{0,1,2}$ and $d^{ijkl}_{0,1,2}$ can be analytically inferred, whose explicit expressions are given in the supplementary materials Supplemental:Material .
ijkl{ijkl} $c^{ijkl}_{0}$ $c^{ijkl}_{2}$ $d^{ijkl}_{0}$ $d^{ijkl}_{2}$ $c^{ijkl}_{1}=d^{ijkl}_{1}$
$0000$ 1.0000 21.334 2.0000 17.491 4.8333
$1000$ 8.1667 24.347 -8.1667 24.347 5.3889
$1010$ 168.78 28.623 174.22 28.251 5.9444
$1100$ 3.5000 28.699 -3.5000 18.233 5.5000
$1110$ 38.111 29.655 -38.111 29.655 6.0556
$1111$ 59.889 27.016 21.778 26.056 6.1667
$2000$ 21.350 26.260 8.0500 24.454 5.7222
$2010$ 89.833 31.367 -89.833 31.367 6.2778
$2011$ 93.100 32.158 9.8000 37.615 6.3889
$2020$ 234.71 34.390 54.880 33.494 6.6111
$2100$ 8.7500 27.639 -8.7500 27.639 5.9444
$2110$ 159.25 32.756 159.25 32.904 6.5000
$2111$ 44.917 32.687 -44.917 32.687 6.6111
$2120$ 84.525 35.895 -84.525 35.895 6.8333
$2121$ 159.25 37.626 159.25 37.603 7.0556
$2200$ 3.2000 27.381 0.10000 45.283 6.0000
$2210$ 12.250 31.624 -12.250 31.624 6.5556
$2211$ 7.1167 39.349 2.6833 35.342 6.6667
$2220$ 30.345 35.977 4.9350 36.841 6.8889
$2221$ 7.1750 38.780 -7.1750 38.780 7.1111
$2222$ 7.7900 34.975 1.7200 34.974 7.1667
@@
We conduct the four-fold convolution integration following two different routes: one is to utilize the numerical integrator Cubature cubature by implementing multiple-precision floating-point around 100 digits, and the other is to carry out the integration analytically with the assistance of the package PolyLogTools Duhr:2019tlz . The perfect agreement is found between the analytic and numerical approaches. In Table 1 we enumerate the values for $c_{i}$ and $d_{i}$ associated with some low-lying $(ijkl)$  [^8] We notice that, a PhD thesis by Knödlseder in 2015 was dedicated to calculating the NLO QCD correction to nucleon Dirac form factor at leading twist Knodlseder:2015vmu . Though the analytic result of the one-loop amplitude was obtained in Knodlseder:2015vmu , the IR-finite hard-scattering kernel $T^{(1)}(x,y)$ has not been given in a pointwise manner. More importantly, since the author failed to accomplish the four-fold convolution integrations in a numerically stable fashion, no attempt of confronting the NLO pQCD predictions with the measured nucleon form factor data was made in Knodlseder:2015vmu ..
Equipped with Table 1, Equation (21) constitutes the master formula of rendering predictions for proton Dirac form factor through one-loop accuracy. In the isospin symmetric limit, the neutron Dirac form factor $F_{1}^{n}$ can be obtained from (21) by making the replacement $e_{u}\leftrightarrow e_{d}$ in (13). It is also straightforward to adapt this master formula from the space-like domain to the time-like one, provided that one makes the replacement $L_{\mu}\to L_{\mu}+i\pi$ in (13), with $Q^{2}$ now indicating the squared invariant mass of the nucleon-antinucleon pair.
Phenomenological exploration. It is now the time to confront the state-of-the-art pQCD predictions with the available large-$Q^{2}$ data. We take the data points for spacelike proton Dirac form factor from Coward:1967au ; Sill:1992qw   [^9] Note it is the magnetic Sachs form factor of nucleon, $G_{\rm M}(Q^{2})$, that is directly accessible in $ep$ scattering experiments. At large $Q^{2}$, it nevertheless becomes legitimate to assume $F_{1}(Q^{2})\approx G_{\rm M}(Q^{2})$., and the data points for time-like proton Dirac form factor from BES:2005lpy ; CLEO:2005tiu ; Seth:2012nn ; E835:1999mlt ; E760:1992rvj ; BaBar:2013ves ; BaBar:2013ukx . We discard many irrelevant data at small momentum transfer. In contrast, the quality of data and the covered range of $Q^{2}$ for neutron EMFFs is much less competitive with respect to that for proton. The data points for spacelike neutron Dirac form factor are taken from Rock:1982gf ; Lung:1992bu , while those for timelike neutron Dirac form factor are taken from Antonelli:1998fv ; BESIII:2021tbq .
@@
As the key nonperturative input, our present knowledge about the nucleon LCDAs is far from confirmative. During the past decades, the low-lying expansion coefficients of nucleon LCDAs have been extensively investigated by different approaches Chernyak:1984bm ; Chernyak:1987nv ; King:1986wi ; Stefanis:1992nw ; Braun:2006hz ; Anikin:2013aka ; Braun:2014wpa ; RQCD:2019hps . In our numerical analysis, we have taken eight different ansatzs of nucleon LCDAs as input. Six of them are obtained from the QCD sum rules, referred to as CZ Chernyak:1984bm , COZ Chernyak:1987nu , KS King:1986wi , SB Stefanis:1992nw , BLW Braun:2006hz and ABO Anikin:2013aka , respectively. Two of them are provided by the recent lattice QCD simulations, abbreviated as LAT 14 Braun:2014wpa and RQCD RQCD:2019hps , respectively.
The predicted values of the expansion coefficients of nucleon are usually given at the initial scale $\mu_{0}=1$ or 2 GeV, which are scattered in a wide range in literature. The one-loop renormalization group equations are invoked to evolve $f_{\rm N}$ and $a_{ij}(\mu)$ from $\mu_{0}$ to another scale, viz., $f_{\rm N}(\mu)=f_{\rm N}(\mu_{0})\left[\alpha_{s}(\mu)/\alpha_{s}(\mu_{0})\right]^{2/(3\beta_{0})}$ and $a_{ij}(\mu)=a_{ij}(\mu_{0})\left[\alpha_{s}(\mu)/\alpha_{s}(\mu_{0})\right]^{\gamma_{ij}/\beta_{0}}$. $\beta_{0}=11-{2\over 3}n_{f}$ with $n_{f}=3$ signifies the one-loop coefficient of the QCD $\beta$ function, while $\gamma_{ij}$ is the anomalous dimension associated with each expansion coefficient, whose explicit form can be found in Braun:2014wpa . The package FAPT Bakulev:2012sm is utilized to evaluate the running QCD coupling constant to two-loop accuracy.
@@
It has been widely known that the LO predictions in pQCD for proton Dirac form factor are generally much smaller than the large-$Q^{2}$ data, for virtually all the popular models of nucleon LCDAs Stefanis:1997zyh  [^10] To resolve this discrepancy, historically it has been proposed Chernyak:1984bm ; Gari:1986ue ; Stefanis:1987vr ; Gari:1986dr ; Ji:1986uh ; Stefanis:1992nw that the strong coupling constants should be evaluated at the exact gluon virtualities in each diagram, so that the LO prediction can be significantly enhanced to account for the data. The $\alpha_{s}$ in this recipe might be associated with a scale of order $\Lambda_{\rm QCD}$, evaluation of which requires some ad hoc treatment. Moreover, it is difficult to adapt this recipe to the NLO case.. As is evident from Fig. 2 through Fig. 5, the magnitudes of the NLO QCD correction are positive and significant. Once we adopt the CZ Chernyak:1984bm , COZ Chernyak:1987nu , KS King:1986wi , SB Stefanis:1992nw parametrizations of nucleon LCDAs, the pQCD predictions at NLO accuracy appear to be in magnitudes compatible with the large-$Q^{2}$ data for both proton and neutron’s Dirac form factors, in both spacelike and timelike regions. Nevertheless, we must stress that the monotonically decreasing tendency of the NLO predictions in variation with $Q^{2}$ is clearly at odds with that exhibited in the actual data.
@@
In sharp contrast, due to quite small expansion coefficients predicted by BLW Braun:2006hz , ABO Anikin:2013aka , LAT14 Braun:2014wpa and RQCD RQCD:2019hps , the resulting pQCD predictions fall far below the experimental data. From Fig. 2 through Fig. 5, one clearly sees that, even after including NLO corrections, the pQCD predictions based on these parametrizations are still order-of-magnitude smaller than the measured Dirac form factor of proton and neutron.
@@
Since lattice QCD is universally believed to be the only model-independent nonperturbative approach to unravel the hadronic structure, the sheer failure of the parameterized nucleon LCDAs provided by LAT14 and RQCD when confronting with the data is rather disquieting. To settle this alarming discrepancy, it seems imperative to call for another independent lattice study based on different algorithm. The large momentum effective theory (LaMET) Ji:2013dva ; Ji:2020ect is a viable option, which allows one to ascertain hadron’s parton distributions directly in the $x$ space on Euclidean lattice Some theoretical exploration to infer light octet baryons’ LCDAs in LaMET approach has recently been launched Deng:2023csv ; Han:2024ucv . Had the prospective LaMET predictions for nucleon LCDAs turned out to be compatible with the RQCD predictions, one would be then inevitably led to the conclusion that the hard-scattering mechanism alone in (2) would be inadequate even at $Q^{2}\approx 30\;{\rm GeV^{2}}$, and some additional soft nonfactorizable contributions are expected to play an important role.
@@
Summary and Outlook. In this work we report the first complete calculation of the NLO QCD corrections to the nucleon Dirac form factors. The collinear factorization formalism has been explicitly verified to hold true at one-loop order for hard-exclusive process involving baryons.
It has long been known that the LO predictions in pQCD severely underestimate the measured nucleon Dirac form factor data. Historically some ad hoc frozen strong coupling recipe were advocated to reconcile this discrepancy. Remarkably, the effect of NLO perturbative corrections turn to be positive and significant. Nevertheless, even after including the sizable NLO QCD corrections, the pQCD predictions based on the recent lattice QCD inputs of nucleon LCDAs, are still considerably smaller than the measured nucleon form factors. This symptom may indicate that the hard-scattering mechanism alone might be inadequate even at $Q^{2}$ accessible in future EIC/EicC experiments, and other soft nonfactorizable contributions need be supplemented to account for the data.
@@
Acknowledgments. We are grateful to Deshan Yang for valuable discussion. L. B. C. is in part supported by National Natural Science Foundation of China (NSFC) under Grant No. 12175048 and the Guangdong Basic and Applied Basic Research Foundation under grant No. 2025B1515020009. W. C. is supported by NSFC under Grant No. 12405095 and Guangdong Major Project of Basic and Applied Basic Research (No. 2020B0301030008). The work of F. F. is supported by NSFC under Grant No. 12275353 and in part by the Fundamental Research Funds for the Central Universities (2025 Basic Sciences Initiative in Mathematics and Physics). The work of S. H. and Y. J. is supported in part by NSFC under Grants No. 12475090 and No. 11925506.
Note added. After the first version of this paper was submitted to arXiv, there has appeared a similar work by Huang, Shi, Wang and Zhao Huang:2024ugd , who have also computed the NLO QCD correction to nucleon Dirac form factor in the spacelike domain, yet within a specific evanescent-to-physical operator mixing scheme. After converted into KM scheme, their results in the supplementary material in the published version of Huang:2024ugd are compatible with ours.
Appendix A Supplementary Material
A.1 One-loop renormalization factor of non-local three-quark operators in KM scheme
In KM scheme Krankl:2011gch , The $\overline{\rm MS}$}-renormalized three-quark operators is connected with the bare one through
where the $\cal Z$} factor can be parameterized as
${\cal Z}_{\alpha\beta\gamma}^{\alpha^{\prime}\beta^{\prime}\gamma^{\prime}}(x,x^{\prime})$ $=$ $\delta_{\alpha\beta\gamma}^{\alpha^{\prime}\beta^{\prime}\gamma^{\prime}}\delta(x-x^{\prime})+\sum_{s=1}{1\over\epsilon^{s}}\,a^{s}_{lmn}\,\gamma^{(l)}_{\alpha\alpha^{\prime}}\otimes\gamma^{(m)}_{\beta\beta^{\prime}}\otimes\gamma^{(n)}_{\gamma\gamma^{\prime}}$ (15)
where $\gamma^{(n)}$ is the shorthand for the totally anti-symmetrized product of $n$ $\gamma$-matrices, $\gamma^{(n)}_{\mu_{1}\cdots\mu_{n}}\equiv\gamma^{[\mu_{1}}\gamma^{\mu_{2}}\cdots\gamma^{\mu_{n}]}$. Note that all the Lorentz indices of $\gamma$-matrices are contracted between different quark lines, where there is only one independent way to contract all indices in $\cal Z$}.
To the best of our knowledge, the explicit one-loop expression for the renormalization factor $\cal Z$} in Eq.(5), which is associated with the nonlocal three-quark operators with light-like separations, has never been explicitly given in literature. For reader’s convenience, here we present its closed one-loop form: